# JailGuard: A Universal Detection Framework for LLM Prompt-based Attacks

### v3

@@ Abstract.
Large Language Models (LLMs) and Multi-Modal LLMs (MLLMs) have played a critical role in numerous applications. However, current LLMs are vulnerable to prompt-based attacks, with jailbreaking attacks enabling LLMs to generate harmful content, while hijacking attacks manipulate the model to perform unintended tasks, underscoring the necessity for detection methods. Unfortunately, existing detecting approaches are usually tailored to specific attacks, resulting in poor generalization in detecting various attacks across different modalities. To address it, we propose JailGuard, a universal detection framework for jailbreaking and hijacking attacks across LLMs and MLLMs. JailGuard operates on the principle that attacks are inherently less robust than benign ones, regardless of method or modality. Specifically, JailGuard mutates untrusted inputs to generate variants and leverages discrepancy of the variants’ responses on the model to distinguish attack samples from benign samples. We implement 18 mutators for text and image inputs and design a mutator combination policy to further improve detection generalization. To evaluate the effectiveness of JailGuard, we build the first comprehensive multi-modal attack dataset, containing 11,000 data items across 15 known attack types. The evaluation suggests that JailGuard achieves the best detection accuracy of 86.14%/82.90% on text and image inputs, outperforming state-of-the-art methods by 11.81%-25.73% and 12.20%-21.40%.
@@
ccs: Security and privacy Software and application securityccs: Computing methodologies Neural networks
@@ 1. Introduction
Large Language Models (LLMs) have become commonplace in our technological interactions, from chatbots to complex decision-making engines (gpt, 2023a; Chen et al., 2021). They can perform various tasks such as understanding sentences, answering questions, etc., and are widely used in many different areas. For example, Meta has developed an AI assistant based on the LLM Llama and integrated it into multiple social platforms such as Facebook (Meta, 2024). The advent of Multi-Modal Large Language Models (MLLMs) has expanded these functionalities even further by incorporating visual understanding, allowing them to interpret and generate imagery alongside text, enhancing user experience with rich, multi-faceted interactions (gpt, 2023b; Zhu et al., 2023a; Liu et al., 2024a). Recently, Microsoft has released Copilot, a search engine based on MLLMs, which supports text and image modal input and provides high-quality information traditional search engines cannot provide (Microsoft, 2024).
@@
Despite the utility of LLMs and MLLMs in various real-world scenarios, they suffer from both jailbreaking and hijacking attacks from different modalities. Jailbreaking attacks can bypass the security mechanisms of the model and software (e.g., AI-driven search engines) and produce harmful or illegal content like sex, violence, abuse, etc (Zou et al., 2023; Chao et al., 2023). Hijacking attacks can hijack and manipulate the LLM systems (e.g., AI assistants) to perform unintended tasks (Liu et al., 2023c). An LLM system might suffer from the two types of attacks on different modalities. For example, an AI assistant that supports multi-modal inputs could be misled by attackers to generate illegal content, or be hijacked to perform unintended tasks, ultimately exposing sensitive information, enabling the spread of misinformation, and damaging the overall trust in AI-driven applications and systems. Thus, there is an urgent need for a universal detection across different attacks and modalities, not only to help prevent these LLM prompt-based attacks and address such security gaps, but the attack samples identified and collected can also help developers understand the attacks and further improve LLM systems and applications.
@@
There are approaches proposed to detect harmful content and the attack intent in models’ inputs and responses (azu, 2023; Jain et al., 2023; Robey et al., 2023). Despite these commendable efforts, existing LLM detection approaches still have limitations, resulting in poor adaptability and generalization across different modalities and attack methods. Typically, these methods rely on specific detection techniques or metrics (e.g., keywords and rules) to identify a limited range of attacks. They are designed to detect either jailbreaking attacks that produce harmful content (azu, 2023) or hijacking attacks that manipulate LLMs to perform specific tasks (Liu et al., 2023c). While such designs perform well on samples generated by specific attack methods, they struggle to detect attacks generated by other methods. Moreover, simply combining these detectors can result in a significant number of false positives. Consequently, existing detection methods are impractical for deployment in real-world scenarios where diverse attacks span different modalities.
@@
To break through these limitations, we design and implement a universal detection framework for LLM prompt-based attacks, JailGuard, which effectively identifies various prompt-based attacks on both image and text modalities. The key observation behind JailGuard is that attack inputs inherently exhibit lower robustness than benign queries, regardless of the attack methods and modalities. To confuse LLMs, attack inputs are often generated based on crafted templates or by an extensive searching process with complex perturbations. This leads to a result that any minor modification to the inputs may invalidate the attack’s effectiveness, which appears as a significant change in output. The responses of benign inputs, however, are hardly affected by these perturbations, as shown in the Fig. 1. Based on this observation, JailGuard first mutates the original input into a series of variant queries. Then the consistency of the responses of LLM systems to variants is analyzed. If a notable discrepancy can be identified among the responses, i.e., a divergence value that exceeds the built-in threshold, a potential prompt-based attack is identified. To effectively identify various attacks, JailGuard systematically designs and implements 16 random mutators and 2 semantic-driven targeted mutators to introduce perturbations at the different levels of text and image inputs. We observe that the detection effectiveness of JailGuard is closely tied to the mutation strategy, as different mutators apply disturbances at various levels and are suitable for detecting different attack methods. To design a more general and effective mutation strategy in detecting a wide range of attacks, JailGuard proposes a mutator combination policy as the default mutation strategy. Based on the empirical data of mutators, the policy selects three mutators to apply perturbations from different levels, combines their variants and divergences according to an optimized probability, and leverages their strengths to detect various attacks comprehensively.
@@
To comprehensively evaluate JailGuard, we construct the first comprehensive LLM prompt-based attack dataset that contains 11,000 items of data covering 15 types of jailbreaking and hijacking attacks on image and text modalities. Based on this dataset, we conduct large-scale experiments that spend over 500M paid tokens to compare JailGuard with 12 state-of-the-art (SOTA) jailbreaking and hijacking detection methods on text and image inputs, including commercial detector Azure content detector (azu, 2023). The experimental results indicate that all mutators in JailGuard can effectively identify LLM prompt-based attacks and benign samples on image and text modalities, achieving higher detection accuracy than SOTA. In addition, the default combination policy of JailGuard further improves the detection results and has separately achieved the best accuracy of 86.14% and 82.90% on text and image inputs, significantly outperforming state-of-the-art defense methods by 11.81%-25.73% and 12.20%-21.40%. In addition, JailGuard can effectively detect and defend different types of prompt-based attacks. Among all types of collected attacks, the best detection accuracy in JailGuard ranges from 76.56% to 100.00%. The default combination policy in JailGuard can achieve an accuracy of more than 70% on 10 types of text attacks, and the detection accuracy on benign samples is over 80%, which exhibits the best generalization among all mutators and baselines. Furthermore, the experiment results also demonstrate the efficiency of JailGuard. We observe that the detection accuracy of the JailGuard’s mutators does not drop significantly when the LLM query budgets (i.e., the number of generated variants) reduce from $N=8$ to $N=4$ and is always better than that of the best baseline SmoothLLM. This finding can provide guidance on attack detection and defense in low-budget scenarios.
In summary, our contributions are:
@@
  • We identify the inherent low robustness of LLM attack inputs. Based on that, we design and implement the first universal LLM attack detection framework, JailGuard, which implements 16 random mutators, 2 semantic-driven targeted mutators, and a set of combination policies. JailGuard mutates the original query to generate variants and uses the divergence of the variants’ responses to detect LLM prompt-based attacks (i.e., jailbreaking and hijacking attacks) on image and text modalities.
@@
  • We construct the first comprehensive LLM attack dataset that consists of 11,000 samples and covers 15 jailbreaking and hijacking attacks on both image and text inputs.
@@
Threat to Validity. JailGuard is currently evaluated on a dataset consisting of 11,000 items of data and 12 attack methods, which may be limited. Although our basic idea can theoretically be applied to other attack methods, this may not hold on some unseen attacks. To mitigate the threats and follow the Open Science Policy, the code of the prototype JailGuard and all the experiment results will be publicly available at (our, 2024).
@@ 2.1. Prompt-based LLM Attack
Prompt-based LLM attack aims to design and generate an attack prompt $P_{a}$ that can induce the target LLM system and application $M$ to contain attack target $T$ in the model response $R=M(P_{a})$, which can be expressed as follows.
@@
where $eval(\cdot)$ is an evaluation function and it returns 1 iff the input prompt $P_{a}$ bypasses the LLM’s protection mechanism and the corresponding response $R$ achieves the attack target $T$. This paper aims to design a detection framework to identify prompt-based attacks, which are mainly divided into jailbreaking and hijacking attacks according to the differences in attack target $T$.
Jailbreaking attack leverages elaborate templates, specific strings, etc. to guide the LLMs to generate toxic contents that violate usage policies [^1] https://platform.openai.com/docs/guides/moderation, such as sexual information and hateful contents To effectively and automatically generate jailbreak prompts, researchers proposed a variety of attack methods (Zou et al., 2023; Deng et al., 2023a; Chao et al., 2023; Zhu et al., 2023b; Yu et al., 2023). Zou et al. (Zou et al., 2023) designed the greedy coordinate gradient-based search (GCG) to produce adversarial suffix to attack open-sourced LLMs (e.g., Vicuna (Zheng et al., 2023)), which has proven its effectiveness through transfer attacks on black-box commercial LLMs. TAP (Mehrotra et al., 2023) is one of the state-of-the-art jailbreaking methods that only requires black-box access to the target LLM. It utilizes LLMs to iteratively refine candidate attack prompts using tree-of-thoughts reasoning until one of the generated prompts jailbreaks the target LLM. With the emergence of MLLMs, researchers designed a visual jailbreaking attack based on the implanted adversarial perturbation in the image inputs (Qi et al., 2024). Their method achieved a high attack success rate on MiniGPT-4 which is one of the state-of-the-art MLLMs (Zhu et al., 2023a). We collect 10 jailbreaking attacks at the text and image level in our dataset, as shown in § 5.
@@ 2.2. LLM Attack Detector
Conducting a detector to identify the attack inputs of the given model is one of the most popular defense strategies in the field of AI security (Metzen et al., 2017; Wang et al., 2021; Liu et al., 2022; Dong et al., 2021). Detectors not only prevent attacks that have not yet occurred on AI systems but also help developers understand attack methods and attacker intentions, thereby improving the safety and security of AI systems and software. Existing LLM attack detectors leverage the model input prompt $P$ and response $R$ to identify attacks. The expected output of the detector can be expressed as follows.
$detect(P,R)=\begin{cases}1&\text{if}\quad P\in\mathcal{P}_{a}\\ 0&\text{otherwise}\end{cases}$
where $\mathcal{P}_{a}$ represents the attack prompt set. When the detector recognizes the attack input, its output is 1 and such an input will be filtered. Otherwise, the output is 0 and the LLM response passes the detector.
@@
LLM attack detectors are widely used in commercial services such as Azure content detector (azu, 2023). Guo et al. (Gou et al., 2024) design the LLM-based harm detector to identify the attack inputs on MLLMs and then regenerate safe-aligned responses. Alon et al. (Alon and Kamfonas, 2023) propose to detect jailbreaking attacks by evaluating the perplexity of queries with adversarial suffixes. Similarly, Liu et al. (Liu et al., 2023c) implement a detection method that uses a built-in threshold and the perplexity of input query and LLMs to identify prompt injection attacks. In addition, SmoothLLM (Robey et al., 2023) uses a set of refusal keywords (Zou et al., 2023) to distinguish blocked jailbreak attack responses from normal responses and aggregates them to obtain the final LLM response. In this paper, we propose a universal LLM attack detection framework, JailGuard. We compare 12 open-sourced LLM detection and defense methods to demonstrate the effectiveness of JailGuard in detecting LLM jailbreaking and hijacking attacks.
3. Motivation

### v4

# JailGuard: A Universal Detection Framework for Prompt-based Attacks on LLM Systems

@@ Abstract.
The systems and software powered by Large Language Models (LLMs) and Multi-Modal LLMs (MLLMs) have played a critical role in numerous scenarios. However, current LLM systems are vulnerable to prompt-based attacks, with jailbreaking attacks enabling the LLM system to generate harmful content, while hijacking attacks manipulate the LLM system to perform attacker-desired tasks, underscoring the necessity for detection tools. Unfortunately, existing detecting approaches are usually tailored to specific attacks, resulting in poor generalization in detecting various attacks across different modalities. To address it, we propose JailGuard, a universal detection framework deployed on top of LLM systems for prompt-based attacks across text and image modalities. JailGuard operates on the principle that attacks are inherently less robust than benign ones. Specifically, JailGuard mutates untrusted inputs to generate variants and leverages the discrepancy of the variants’ responses on the target model to distinguish attack samples from benign samples. We implement 18 mutators for text and image inputs and design a mutator combination policy to further improve detection generalization. The evaluation on the dataset containing 15 known attack types suggests that JailGuard achieves the best detection accuracy of 86.14%/82.90% on text and image inputs, outperforming state-of-the-art methods by 11.81%-25.73% and 12.20%-21.40%.
@@
copyright: acmlicensedjournal: TOSEMjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3724393ccs: Security and privacy Software and application securityccs: Computing methodologies Neural networks
@@ 1. Introduction
In the era of Software Engineering (SE) 3.0, software and systems driven by Large Language Models (LLMs) have become commonplace, from chatbots to complex decision-making engines (gpt, 2023a; Chen et al., 2021; Hassan et al., 2024). They can perform various tasks such as understanding sentences, answering questions, etc., and are widely used in many different areas. For example, Meta has developed an AI assistant based on the LLM ‘Llama’ and integrated it into multiple social platforms such as Facebook (Meta, 2024). The advent of Multi-Modal Large Language Models (MLLMs) has expanded these functionalities even further by incorporating visual understanding, allowing them to interpret and generate imagery alongside text, enhancing user experience with rich, multi-faceted interactions (gpt, 2023b; Zhu et al., 2023a; Liu et al., 2024c). Recently, Microsoft has released Copilot, a search engine based on MLLMs, which supports text and image modal input and provides high-quality information traditional search engines cannot provide (Microsoft, 2024).
@@
As the key component of the LLM system, LLMs are predominantly deployed remotely, requiring users to provide prompts through designated interfaces of systems and software to assess them. While these systems have demonstrated strong utility in various real-world applications, they are vulnerable to prompt-based attacks (e.g., jailbreaking and hijacking attacks) across various modalities. Prompt-based attacks manipulate the output of LLM with carefully designed prompts, thus attacking and endangering the entire system and software. Jailbreaking attacks can circumvent the built-in safety mechanisms of LLM systems (e.g., AI-powered search engines), enabling the systems to generate harmful or illegal content like sex, violence, abuse, etc (Zou et al., 2023; Chao et al., 2023), thereby posing significant security risks. The severity of this security risk is exemplified by a recent incident where a user exploited one of the most popular LLM systems, ChatGPT, to plan and carry out bomb attacks (Press, 2025). Hijacking attacks can hijack and manipulate LLM systems (e.g., AI assistants) to perform specific tasks and return attacker-desired results to the user, thereby disabling the LLM system or performing unintended tasks, jeopardizing user interests and safety. For example, hijacking attacks can manipulate an LLM-based automated screening application to directly generate a response of ‘Hire him’ for the target resume, regardless of its content (Liu et al., 2024b). An LLM system might suffer from the two types of attacks on different modalities. For example, an AI assistant that supports multi-modal inputs could be misled by attackers to generate illegal content, or be hijacked to perform unintended tasks and return attacker-desired results, ultimately exposing sensitive information, enabling the spread of misinformation, and damaging the overall trust in AI-driven software and systems. Thus, there is an urgent need to design and implement universal detection for prompt-based attacks on LLM systems and software, not only to help prevent these attacks across different modalities and address such security gaps, but attack samples identified and collected can also help developers understand the attacks and further improve LLM systems and software.
@@
There are approaches proposed to detect attacks based on models’ inputs and responses (azu, 2023; Jain et al., 2023; Robey et al., 2023). Despite these commendable efforts, existing LLM attack detection approaches still have limitations, resulting in poor adaptability and generalization across different modalities and attack methods. Typically, these methods rely on specific detection techniques or metrics (e.g., keywords and rules) to identify a limited range of attacks. They are designed to detect either jailbreaking attacks that produce harmful content (azu, 2023) or hijacking attacks that manipulate LLMs to generate attacker-desired content (Liu et al., 2024b). While such designs perform well on samples generated by specific attack methods, they struggle to detect attacks generated by other methods. Moreover, simply combining these detectors can result in a significant number of false positives in attack detection. Consequently, existing detection methods are impractical for deployment in real-world LLM systems facing diverse attacks spanning different modalities.
To break through these limitations, we design and implement a universal detection framework for the prompt-based attacks on LLM systems, JailGuard. Developers can deploy JailGuard on the top of the LLM systems as a detection module which can effectively identify various prompt-based attacks on both image and text modalities. The key observation behind JailGuard is that attack inputs inherently exhibit lower robustness on textual features than benign queries, regardless of the attack methods and modalities. For example, in the case of text inputs, when subjected to token or word level perturbations that do not alter the overall semantics, attack inputs are less robust than benign inputs and are prone to failure. The root cause is that to confuse the model in LLM systems, attack inputs are often generated based on crafted templates or by an extensive searching process with complex perturbations. This results in any minor modification to the inputs that may invalidate the attack’s effectiveness, which manifests as a significant change in output and a large divergence between the LLM responses. The responses of benign inputs, however, are hardly affected by these perturbations. Fig. 1 provides a demo case of this observation. We use heat maps to intuitively show the divergence of the LLM responses to benign inputs and the divergence of the responses to attack inputs. Compared to benign inputs, variants of attack inputs can lead to greater divergences between LLM responses, which can be used to identify attack inputs. Based on this observation, JailGuard first mutates the original input into a series of variant queries. Then the consistency of the responses of LLMs to variants is analyzed. If a notable discrepancy can be identified among the responses, i.e., a divergence value that exceeds the built-in threshold, a potential prompt-based attack is identified. To effectively identify various attacks, JailGuard systematically designs and implements 16 random mutators and 2 semantic-driven targeted mutators to introduce perturbations at the different levels of text and image inputs. We observe that the detection effectiveness of JailGuard is closely tied to the mutation strategy, as different mutators apply disturbances at various levels and are suitable for detecting different attack methods. To design a more general and effective mutation strategy in detecting a wide range of attacks, JailGuard proposes a mutator combination policy as the default mutation strategy. Based on the empirical data of mutators on the development set, the policy selects three mutators to apply perturbations from different levels, combines their variants and divergences according to an optimized probability, and leverages their strengths to detect various attacks comprehensively.
@@
To evaluate the effectiveness of JailGuard, we construct the first comprehensive prompt-based attack dataset that contains 11,000 items of data covering 15 types of jailbreaking and hijacking attacks on image and text modalities that can successfully attack GPT-3.5-turbo-1106 and MiniGPT-4 models. These models are widely embedded in LLM systems and software (Kernan Freire et al., 2023). Based on this dataset, we conduct large-scale experiments that spend over 500M paid tokens to compare JailGuard with 12 state-of-the-art (SOTA) jailbreaking and hijacking detection methods on text and image inputs, including commercial detector Azure content detector (azu, 2023). The experimental results indicate that all mutators in JailGuard can effectively identify prompt-based attacks and benign samples on image and text modalities, achieving higher detection accuracy than SOTA. In addition, the default combination policy of JailGuard further improves the detection results and has separately achieved the best accuracy of 86.14% and 82.90% on text and image inputs, significantly outperforming state-of-the-art defense methods by 11.81%-25.73% and 12.20%-21.40%. In addition, JailGuard can effectively detect and defend different types of prompt-based attacks. Among all types of collected attacks, the best detection accuracy in JailGuard ranges from 76.56% to 100.00%. The default combination policy in JailGuard can achieve an accuracy of more than 70% on 10 types of text attacks, and the detection accuracy on benign samples is over 80%, which exhibits the best generalization among all mutators and baselines. Furthermore, the experiment results also demonstrate the efficiency of JailGuard. We observe that the detection accuracy of the JailGuard’s mutators does not drop significantly when the LLM query budgets (i.e., the number of generated variants) reduce from $N=8$ to $N=4$ and is always better than that of the best baseline SmoothLLM. This finding can provide guidance on attack detection and defense in low-budget scenarios. In summary, our contributions are:
@@
  • We identify the inherent low robustness of prompt-based attacks on LLM systems. Based on that, we design and implement the first universal prompt-based attack detection framework, JailGuard, which implements 16 random mutators, 2 semantic-driven targeted mutators, and a set of combination policies. JailGuard can be deployed on the top of LLM systems and it mutates the model input in the LLM system to generate variants and uses the divergence of the variants’ responses to detect the prompt-based attacks (i.e., jailbreaking and hijacking attacks) on image and text modalities.
@@
  • We construct the first comprehensive prompt-based attack dataset that consists of 11,000 samples and covers 15 jailbreaking and hijacking attacks on both image and text inputs, aiming to promote future security research on LLM systems and software.
@@
Threat to Validity. JailGuard is currently evaluated on a dataset consisting of 11,000 items of data and 15 attack methods, which may be limited. Although our basic idea can theoretically be extended to detect other attack methods, this may still fail on some unseen attacks. Moreover, the hyperparameters are model-specific in JailGuard and are obtained through large-scale evaluation of thousands of items of data with 15 attack methods. Although they have achieved excellent detection results in experiments, the detection performance may not be maintained on unseen attacks. We recommend that users tune the hyperparameters (e.g., the selected mutators and probabilities in the combination policy) based on their target LLM system before deployment to achieve optimal performance. To mitigate the threats and follow the Open Science Policy, the code of the prototype JailGuard and all the experiment results will be publicly available at (our, 2024).
@@ 2. Background
2.1. LLM System
LLM-powered systems have emerged as a variety of tools capable of performing diverse tasks, including question-answering, reasoning, and code generation  (Weber, 2024; Parnin et al., 2023). These LLM systems receive and process queries from users, complete downstream tasks embedded in their design and finally return the task results as the system output, such as reasoning answers and the generated code. These systems operate through a three-stage pipeline, namely processing input, querying LLM, and executing downstream task, as illustrated in Fig. 2.
Processing input receives and transforms user input into system-specific model inputs. This transformation varies based on the system’s design and application context, potentially incorporating templates or supplementary details (Kernan Freire et al., 2023; Cai et al., 2023; dal, 2023). To ensure precise query execution, many systems provide users with direct access to write and edit the model input (Cai et al., 2023).
Querying LLM represents the system’s core functionality. In this stage, the processed inputs are submitted to the target LLM (i.e., the key component of the LLM system) to generate responses, as shown in the dashed box in Fig. 2. Since attackers typically lack access to remotely deployed models, they often resort to various prompt-based attacks, crafting specialized inputs to manipulate model responses. To address this security concern, we design and implement JailGuard, a universal detection framework for prompt-based attacks on different modalities, which is deployed in the LLM system and operated before the querying stage.
Executing downstream task leverages specialized software and tools to process LLM responses for specific applications (Kumar et al., 2023b; Mees et al., 2023). For example, in the code generation and question-answering scenario, this stage involves formatting and visually presenting the generated code or answers to users (Cai et al., 2023; Parnin et al., 2023). Similarly, in the scenario of automated screening in hiring, the system can automatically dispatch emails to administrators or applicants based on LLM responses.
Fig. 2. The Basic Stages of the LLM System
@@ 2.2. Prompt-based LLM Attack
Existing LLMs are usually safety-aligned and often provide refusal responses to the straightforward harmful prompts (e.g., ‘how to make bombs’) queried by attackers (Zou et al., 2023; Cui et al., 2024). However, the safety alignment mechanism of LLM can not block all harmful prompts. Prompt-based LLM attack aims to design and generate an attack prompt $P_{a}$ that can bypass the safety alignment and induce the model $M$ in the target system or software to contain attack target $T$ in the model response $R=M(P_{a})$, which can be expressed as follows.
@@
where $eval(\cdot)$ is an evaluation function and it returns 1 iff the input prompt $P_{a}$ bypasses the LLM’s protection mechanism and the corresponding response $R$ achieves the attack target $T$. This paper aims to design a detection framework to identify prompt-based attacks that can obtain 1 in evaluation $eval(\cdot)$. The attacks are mainly divided into jailbreaking and hijacking attacks according to the differences in attack target $T$.
Jailbreaking attack leverages elaborate templates, specific strings, etc. to guide the LLMs to generate toxic contents that violate usage policies(e.g., OpenAI policy [^1] https://platform.openai.com/docs/guides/moderation), such as sexual information and hateful contents. The left part of Fig. 3 provides a demo case of jailbreaking attack (Wei et al., 2024) on GPT-3.5-1106 model. The input prompts constructed by the attackers successfully bypass the LLM system’s safety alignment, leading the model to generate harmful content about how to promote and market adult services effectively. For LLM systems and software in the question-answering scenario, such a harmful response will be returned and displayed to users, which violates the usage policies. To effectively and automatically generate jailbreak prompts, researchers proposed a variety of attack methods (Zou et al., 2023; Deng et al., 2023a; Chao et al., 2023; Zhu et al., 2023b; Yu et al., 2023). Zou et al. (Zou et al., 2023) design the greedy coordinate gradient-based search (GCG) to produce adversarial suffix to attack open-sourced LLMs (e.g., Vicuna (Zheng et al., 2023)), which has proven its effectiveness through transfer attacks on black-box commercial LLMs. TAP (Mehrotra et al., 2023) is one of the state-of-the-art jailbreaking methods that only requires black-box access to the target LLM. It utilizes LLMs to iteratively refine candidate attack prompts using tree-of-thoughts reasoning until one of the generated prompts jailbreaks the target LLM. With the emergence of MLLMs, researchers design visual jailbreaking attacks by implanting adversarial perturbation in the image inputs (Qi et al., 2024). Their method achieved a high attack success rate on MiniGPT-4 which is one of the state-of-the-art MLLMs (Zhu et al., 2023a). We collect a total of 8 jailbreaking attacks at the text and image level in our dataset, as shown in Table 2.
Fig. 3. Demo Cases of Jailbreaking and Hijacking Attacks
@@ 2.3. LLM Attack Detector
Conducting a detector to identify the attack inputs of the given model is one of the most popular defense strategies (Metzen et al., 2017; Wang et al., 2021a; Liu et al., 2022; Dong et al., 2021). Detectors not only prevent attacks that can bypass the safety mechanism of LLM systems but also help developers understand attack methods and attacker intentions, thereby improving the safety and security of LLM systems and software. For instance, after the detector identifies and blocks the attack prompts, it can save them and build a real-world attack dataset. On the one hand, developers can analyze the templates and methods of these attack prompts, study the attack target and intentions, and further design and implement targeted defense mechanisms for LLM systems (Xie et al., 2023; Robey et al., 2023). On the other hand, they can directly leverage the collected attack dataset to conduct continuous learning and safety alignment (Shen et al., 2024a; Touvron et al., 2023) on LLMs, inherently improving the safety and security of the LLM system and software. Therefore, designing attack detectors to identify prompt-based attacks on LLM systems is of great importance for improving software quality and security in the era of LLMs (Liu et al., 2024b, a; Hassan et al., 2024; Martínez-Fernández et al., 2022).
Existing LLM attack detectors leverage the model input prompt $P$ and response $R$ to identify attacks. The expected output of the detector can be expressed as follows.
(2) $detect(P,R)=\begin{cases}1&\text{if}\quad P\in\mathcal{P}_{a},\\ 0&\text{otherwise},\end{cases}$
where $\mathcal{P}_{a}$ represents the attack prompt set. When the detector recognizes the attack input, its output is 1 and such an input will be filtered. Otherwise, the output is 0 and the LLM response passes the detector. Note that the LLM attack detector is usually implemented on the top of the LLM systems to prevent the system from prompt-based attacks. The attack prompt set $\mathcal{P}_{a}$ it detects consists of valid attack prompts that can lead to successful attacks (e.g., guiding the model $M$ to generate harmful contents). Those samples that fail to achieve attacks on the LLM system have little significance for developers and the security of LLM systems, and they cannot reflect potential problems and defects in the LLM system. To ensure the data quality, all attack samples collected in § 5 and used in our experiments have been verified to be able to successfully attack the model $M$ in the target system.
@@
To effectively detect these valid attack prompts, researchers have proposed various methods, which can be divided into the pre-query method and the post-query method. Post-query methods detect the LLM attacks after the querying LLM stage in Fig. 2. Commercial content detectors (e.g., Azure content detector (azu, 2023)) commonly used in LLM systems usually belong to this category. They leverage the model’s responses to the original prompt to determine whether this input is harmful. Guo et al. (Gou et al., 2024) design the LLM-based harm detector to identify the attack inputs based on MLLM responses to the given inputs and then regenerate safe-aligned responses. Since post-query detectors usually leverage built-in rules, thresholds, and integrated models to identify harmful content, this makes them heavily influenced by the design of the rules and susceptible to false negatives for unknown attacks. Pre-query methods detect attacks before the querying LLM stage. For the pre-query defense, one of the state-of-the-art methods is SmoothLLM (Robey et al., 2023), which mutates the original inputs and uses a set of refusal keywords (Zou et al., 2023) to distinguish blocked jailbreak attack responses from normal responses and aggregates them to obtain the final LLM response. Alon et al. (Alon and Kamfonas, 2023) propose to detect jailbreaking attacks by evaluating the perplexity of queries with adversarial suffixes. Similarly, Liu et al. (Liu et al., 2024b) implement a detection method that uses a built-in threshold and the perplexity of input query and LLMs to identify prompt injection attacks. Regardless of how the detection methods are designed, both pre-query and post-query methods share the same task, to detect and prevent prompts that can attack the LLM. In this paper, we propose a universal LLM attack detection framework for such a task, JailGuard. Sharing the same position as popular commercial detectors used in LLM systems (e.g., Azure content detector), JailGuard aims to detect and identify various prompt-based attacks that can attack and harm LLM systems. We compare 12 open-sourced LLM pre-query and post-query detection and defense methods to demonstrate the effectiveness of JailGuard in detecting LLM jailbreaking and hijacking attacks.
2.4. Kullback-Leibler Divergence
Kullback-Leibler (KL) divergence measures the difference between two probability distributions and is widely used in NLP tasks (Zhu et al., 2019; Hare et al., 2020; Huang et al., 2008; Kumar et al., 2023c). It can be formulated as follows.
(3) $D(P\parallel Q)=\sum_{x}P(x)\log\left(\frac{P(x)}{Q(x)}\right),$
where $P$ and $Q$ are probability distributions. Since the KL divergence is non-negative, it reaches the minimum when two distributions are the same (i.e., $P=Q$). JailGuard employs KL divergence to quantify the differences between the similarity distributions of LLM responses, effectively identifying attack prompts that are susceptible to perturbations and result in divergent outputs.
3. A closer look at the motivation